# Elucidating the Design Space of Arbitrary-Noise-Based Diffusion Models

### v1

@@ Appendix D Derivation of Optimal Case for EDA
$f(t)=\frac{s^{\prime}(t)}{s(t)},\quad\phi(t)=\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m}h_{m},\quad g(t)=\frac{s(t)}{\eta+1}\sqrt{\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}}$ (80)
Suppose our training set consists of a finite samples $[y_{1},\ldots,y_{Y}]$. Thus $p_{\text{data}}(x)$ is represented by a mixture of Dirac delta distributions:
@@
$p(x)$ $=\int_{\mathbb{R}^{d}}p_{\text{data}}(x_{0})\mathcal{N}\left(x;s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)\mathrm{d}x_{0}$ (82)
$=\int_{\mathbb{R}^{d}}\frac{1}{Y}\sum_{i=1}^{Y}\delta(x_{0}-y_{i})\mathcal{N}\left(x;s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)\mathrm{d}x_{0}$ (83)
$=\frac{1}{Y}\sum_{i=1}^{Y}\int_{\mathbb{R}^{d}}\delta(x_{0}-y_{i})\mathcal{N}\left(x;s(t)x_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)\mathrm{d}x_{0}$ (84)
$=\frac{1}{Y}\sum_{i=1}^{Y}\mathcal{N}\left(x;s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)$ (85)
@@
$\mathcal{L}(D;\sigma)$ $=\mathbb{E}_{y\sim p_{\text{data}}}\mathbb{E}_{x\sim p(x|y)}\|D(x;\sigma)-y\|^{2}$ (86)
$=\mathbb{E}_{y\sim p_{\text{data}}}\int_{\mathbb{R}^{d}}\mathcal{N}\left(s(t)y+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)\|D(x;\sigma)-y\|^{2}\mathrm{d}x$ (87)
$=\frac{1}{Y}\sum_{i=1}^{Y}\int_{\mathbb{R}^{d}}\mathcal{N}\left(s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)\|D(x;\sigma)-y_{i}\|^{2}\mathrm{d}x$ (88)
$=\int_{\mathbb{R}^{d}}\frac{1}{Y}\sum_{i=1}^{Y}\mathcal{N}\left(s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)\|D(x;\sigma)-y_{i}\|^{2}\mathrm{d}x$ (89)
$\triangleq\int_{\mathbb{R}^{d}}\mathcal{L}(D;x,\sigma)\mathrm{d}x$ (90)
We minimize $\mathcal{L}(D;\sigma)$ by minimizing $\mathcal{L}(D;x,\sigma)$ for each individual $x$. The $\mathcal{L}(D;x,\sigma)$ is a convex optimization problem whose solution can be computed by taking the gradient to zero:
00 0 $=\nabla_{D(x,\sigma)}\mathcal{L}(D;x,\sigma)$ (91)
00 0 $=\nabla_{D(x,\sigma)}\frac{1}{Y}\sum_{i=1}^{Y}\mathcal{N}(\cdot)\|D(x;\sigma)-y_{i}\|^{2}\mathrm{d}x$ (92)
00 0 $=\frac{1}{Y}\sum_{i=1}^{Y}\mathcal{N}(\cdot)\nabla_{D(x,\sigma)}\|D(x;\sigma)-y_{i}\|^{2}\mathrm{d}x$ (93)
00 0 $=\sum_{i=1}^{Y}\mathcal{N}(\cdot)\left[2D(x;\sigma)-2y_{i}\right]$ (94)
$D(x;\sigma)$ $=\frac{\sum_{i=1}^{Y}\mathcal{N}(\cdot)y_{i}}{\sum_{i=1}^{Y}\mathcal{N}(\cdot)}$ (95)
where $\mathcal{N}(\cdot)$ denotes the Gaussian probability density function from previous Eq. 89.
@@
$=\frac{\nabla_{x_{t}}\frac{1}{Y}\sum_{i=1}^{Y}\mathcal{N}\left(x_{t};s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)}{\frac{1}{Y}\sum_{i=1}^{Y}\mathcal{N}\left(x_{t};s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)}$ (97)
$=\frac{\sum_{i=1}^{Y}\nabla_{x_{t}}\mathcal{N}\left(x_{t};s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)}{\sum_{i=1}^{Y}\mathcal{N}\left(x_{t};s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)}$ (98)
where the $\nabla_{x_{t}}\mathcal{N}(\cdot)$ is:
$\nabla_{x_{t}}\mathcal{N}\left(x_{t};s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m}h_{m},\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\right)$ (99)
$=\mathcal{N}(\cdot)\nabla_{x_{t}}\left[-\frac{(\eta+1)^{2}}{2s^{2}(t)\sigma^{2}(t)}\left(s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-x_{t}\right)^{\top}\mathbf{\Sigma}^{-1}\right.$
$\qquad\left.\left(s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-x_{t}\right)\right]$ (100)
$=\mathcal{N}(\cdot)\left[\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\mathbf{\Sigma}^{-1}\left(s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-x_{t}\right)\right]$ (101)
Substituting the $\nabla_{x_{t}}\mathcal{N}(\cdot)$ in Eq. 98 yields:
$\nabla_{x_{t}}\log p(x_{t})$ $=\frac{\sum_{i=1}^{Y}\mathcal{N}(\cdot)\left[\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\mathbf{\Sigma}^{-1}\left(s(t)y_{i}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-x_{t}\right)\right]}{\sum_{i=1}^{Y}\mathcal{N}(\cdot)}$ (102)
$=\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\mathbf{\Sigma}^{-1}\left[\frac{s(t)\sum_{i=1}^{Y}\mathcal{N}(\cdot)y_{i}}{\sum_{i=1}^{Y}\mathcal{N}(\cdot)}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-x_{t}\right]$ (103)
Substituting $D(x;\sigma)$ into Eq.103:
$\nabla_{x_{t}}\log p(x_{t})=\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\mathbf{\Sigma}^{-1}\left[s(t)D(x_{t};\sigma)+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-x_{t}\right]$ (104)
Substituting the score function $\nabla_{x_{t}}\log p(x_{t})$ into PFODE, and the final deterministic sampling formula is:
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=\frac{s^{\prime}(t)}{s(t)}x+\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-\frac{s^{2}(t)\sigma(t)\sigma^{\prime}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\left(\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\mathbf{\Sigma}^{-1}\right.$
$\qquad\left.\left(s(t)D(x;\sigma)+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m}-x\right)\right)$ (105)
$=\left(\frac{s^{\prime}(t)}{s(t)}+\frac{\sigma^{\prime}(t)}{\sigma(t)}\right)x-\frac{\sigma^{\prime}(t)s(t)}{\sigma(t)}D(x;\sigma)$
@@ Appendix E Implementation Details
E.1 Bias Field correction in MRI
@@
where the range of both $x$ and $y$ coordinates is $[-1,1]$. Each basis function in $H_{N_{1},N_{2}}(x,y)$ is mapped linearly to the range centered at $[0.9,1.1]$ for even fairness. We use $H_{3,5}(x,y)$ as basis set. The neural network $\theta$ predicts the diffused noise in $x_{t}$ referred to [6, 21], and the loss function is:
@@
E.2 Metal Artifact Reduction in CT
@@
$\begin{split}L_{MAR}(\theta)=&\mathbb{E}\left(1+(1-m)\dfrac{max(m\cdot{N})}{max((1-m)\cdot{N})}\right)\|{N_{\theta}}(x_{t},t)-{N}\|_{2}^{2}\end{split}$ (108)
@@
E.3 Shadow Removal in Natural Image
@@
The diffused noise ${N}$ as $(10+\epsilon)({X_{S}}-{x_{0}})/11$ similar to MAR, where the ${X_{S}}$ is shadow image and ${x_{0}}$ is non-shadow image. Our learning objective should be consistent with the baseline model ShadowFormer that directly learns the non-shadow image ${x_{0}}$. Hence the loss function is:

### v2

@@ Appendix D Derivation of Optimal Case for EDA
$f(t)=\frac{s^{\prime}(t)}{s(t)},\quad\phi(t)=\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m}h_{m},$
$\quad g(t)=\frac{s(t)}{\eta+1}\sqrt{\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}}$ (77)
Suppose our training set consists of a finite samples $y_{1},\ldots,y_{Y}]$. Thus $p_{\text{data}}(x)$ is represented by a mixture of Dirac delta distributions:
@@
$p(x)$ $=\int_{\mathbb{R}^{d}}p_{\text{data}}(x_{0})\,\mathcal{N}\!\bigl(x;\,\mu(x_{0},t),\,\Sigma_{t}\bigr)\,\mathrm{d}x_{0}$ (79)
$=\int_{\mathbb{R}^{d}}\tfrac{1}{Y}\!\sum_{i=1}^{Y}\delta(x_{0}-y_{i})\,\mathcal{N}\!\bigl(x;\,\mu(x_{0},t),\,\Sigma_{t}\bigr)\,\mathrm{d}x_{0}$ (80)
$=\tfrac{1}{Y}\!\sum_{i=1}^{Y}\int_{\mathbb{R}^{d}}\delta(x_{0}-y_{i})\,\mathcal{N}\!\bigl(x;\,\mu(x_{0},t),\,\Sigma_{t}\bigr)\,\mathrm{d}x_{0}$ (81)
$=\tfrac{1}{Y}\!\sum_{i=1}^{Y}\mathcal{N}\!\bigl(x;\,\mu(y_{i},t),\,\Sigma_{t}\bigr),$ (82)
where
$\mu(x_{0},t)$ $=s(t)x_{0}+\tfrac{\eta\,s(t)\,\sigma(t)}{\eta+1}\!\sum_{m}h_{m},\qquad$
$\Sigma_{t}=\tfrac{s^{2}(t)\,\sigma^{2}(t)}{(\eta+1)^{2}}\bm{\Sigma}.$ (83)
@@
$\mathcal{L}(D;\sigma)$ $=\mathbb{E}_{y\sim p_{\text{data}}}\mathbb{E}_{x\sim p(x|y)}\|D(x;\sigma)-y\|^{2},$ (84)
$=\mathbb{E}_{y\sim p_{\text{data}}}\int_{\mathbb{R}^{d}}\mathcal{N}\!\bigl(x;\mu_{y}(t),\bar{\Sigma}(t)\bigr)\|D(x;\sigma)-y\|^{2}\,\mathrm{d}x,$ (85)
$=\tfrac{1}{Y}\!\sum_{i=1}^{Y}\int_{\mathbb{R}^{d}}\mathcal{N}\!\bigl(x;\mu_{y_{i}}(t),\bar{\Sigma}(t)\bigr)\|D(x;\sigma)-y_{i}\|^{2}\,\mathrm{d}x,$ (86)
$=\int_{\mathbb{R}^{d}}\!\tfrac{1}{Y}\!\sum_{i=1}^{Y}\mathcal{N}\!\bigl(x;\mu_{y_{i}}(t),\bar{\Sigma}(t)\bigr)\|D(x;\sigma)-y_{i}\|^{2}\,\mathrm{d}x$ (87)
$\triangleq\int_{\mathbb{R}^{d}}\mathcal{L}(D;x,\sigma)\,\mathrm{d}x.$ (88)
where
$\mu_{y}(t)$ $=s(t)y+\frac{\eta\,s(t)\,\sigma(t)}{\eta+1}\sum_{m}h_{m},$
$\bar{\Sigma}(t)$ $=\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}.$
We minimize $\mathcal{L}(D;\sigma)$ by pointwise minimization: for each fixed $x$, the objective reduces to a weighted least-squares problem
$0$ $=\nabla_{D(x;\sigma)}\mathcal{L}(D;x,\sigma)$ (89)
$=\nabla_{D(x;\sigma)}\tfrac{1}{Y}\sum_{i=1}^{Y}w_{i}(x)\,\|D(x;\sigma)-y_{i}\|^{2}$ (90)
$=\tfrac{2}{Y}\sum_{i=1}^{Y}w_{i}(x)\,\bigl(D(x;\sigma)-y_{i}\bigr),$ (91)
$\Rightarrow\quad D(x;\sigma)$ $=\frac{\sum_{i=1}^{Y}w_{i}(x)\,y_{i}}{\sum_{i=1}^{Y}w_{i}(x)}.$ (92)
where
$w_{i}(x)$ $=\mathcal{N}\!\bigl(x;\mu_{y_{i}}(t),\bar{\Sigma}(t)\bigr),$
$\mu_{y_{i}}(t)$ $=s(t)y_{i}+\tfrac{\eta\,s(t)\,\sigma(t)}{\eta+1}\sum_{m}h_{m},$
$\bar{\Sigma}(t)$ $=\tfrac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}.$
and $D(x;\sigma)$ denotes the pointwise minimizer (denoiser output) at location $x$ and noise level $\sigma$.
@@
$=\frac{\nabla_{x_{t}}\tfrac{1}{Y}\!\sum_{i=1}^{Y}\mathcal{N}\!\bigl(x_{t};\mu_{i}(t),\bar{\Sigma}(t)\bigr)}{\tfrac{1}{Y}\!\sum_{i=1}^{Y}\mathcal{N}\!\bigl(x_{t};\mu_{i}(t),\bar{\Sigma}(t)\bigr)}$ (94)
$=\frac{\sum_{i=1}^{Y}\nabla_{x_{t}}\mathcal{N}\!\bigl(x_{t};\mu_{i}(t),\bar{\Sigma}(t)\bigr)}{\sum_{i=1}^{Y}\mathcal{N}\!\bigl(x_{t};\mu_{i}(t),\bar{\Sigma}(t)\bigr)}.$ (95)
where the term $\nabla_{x_{t}}\mathcal{N}(\cdot)$ is given by:
$\nabla_{x_{t}}\mathcal{N}\!\bigl(x_{t};\mu_{i}(t),\bar{\Sigma}(t)\bigr)=\mathcal{N}\!\bigl(x_{t};\mu_{i}(t),\bar{\Sigma}(t)\bigr)$
$\nabla_{x_{t}}\!\Bigl[-\tfrac{1}{2}(x_{t}-\mu_{i}(t))^{\!\top}\!\bar{\Sigma}^{-1}(x_{t}-\mu_{i}(t))\Bigr]$ (96)
$=\mathcal{N}\!\bigl(x_{t};\mu_{i}(t),\bar{\Sigma}(t)\bigr)\,\bar{\Sigma}^{-1}\!\bigl(\mu_{i}(t)-x_{t}\bigr)$
$\cdot\tfrac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}.$ (97)
Substituting $\nabla_{x_{t}}\mathcal{N}(\cdot)$ into Eq. 95 yields:
$\nabla_{x_{t}}\log p(x_{t})$ $=A(t)\frac{\sum_{i}\mathcal{N}_{i}\,[s(t)y_{i}+B(t)-x_{t}]}{\sum_{i}\mathcal{N}_{i}}$
$=A(t)\!\left[s(t)\frac{\sum_{i}\mathcal{N}_{i}y_{i}}{\sum_{i}\mathcal{N}_{i}}+B(t)-x_{t}\right].$ (98)
Substituting $D(x_{t};\sigma)$ gives:
$\nabla_{x_{t}}\log p(x_{t})$ $=A(t)\,[s(t)D(x_{t};\sigma)+B(t)-x_{t}].$ (99)
Finally, substituting Eq. 99 into the PFODE yields the deterministic sampling form:
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=\frac{s^{\prime}(t)}{s(t)}x+\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\!\sum_{m}h_{m}$
$-s^{2}(t)\sigma(t)\sigma^{\prime}(t)\!\left[\frac{A(t)}{(\eta+1)^{2}}(s(t)D(x;\sigma)+B(t)-x)\right]$
$=\!\left(\frac{s^{\prime}(t)}{s(t)}+\frac{\sigma^{\prime}(t)}{\sigma(t)}\right)x-\frac{\sigma^{\prime}(t)s(t)}{\sigma(t)}D(x;\sigma).$
where for compactness we define:
$\mathcal{N}_{i}$ $=\mathcal{N}\!\Bigl(x_{t};\,s(t)y_{i}+\tfrac{\eta s(t)\sigma(t)}{\eta+1}\!\sum_{m}h_{m},\,\tfrac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\mathbf{\Sigma}\Bigr),$
$A(t)$ $=\tfrac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\mathbf{\Sigma}^{-1},\qquad B(t)=\tfrac{\eta s(t)\sigma(t)}{\eta+1}\!\sum_{m=1}^{M}h_{m}.$
@@ Appendix E Implementation Details
E.1 Specific Experimental Setup
For BFC, the mediator $\eta=0$, the basis set $H$ was set as Case 1 in Prop. 1. The input MRIs were $256\times 256$ and trained for $500$ epochs with a batch size of $1$. For MAR and SR, the $\eta=10$ and the basis set $H$ followed Case 2 in Prop. 1. The input CTs were $416\times 416$ and trained $1000$ epochs with a batch size of $1$. The Shadow Images were $256\times 256\times 3$, trained $2000$ epochs with a batch size of $4$.
E.2 Bias Field correction in MRI
@@
where the range of both $x$ and $y$ coordinates is [1,1][-1,1]. Each basis function in $H_{N_{1},N_{2}}(x,y)$ is mapped linearly to the range centered at $[0.9,1.1]$ for even fairness. We use $H_{3,5}(x,y)$ as basis set. The neural network $\theta$ predicts the diffused noise in $x_{t}$ referred to [9, 28], and the loss function is:
@@
E.3 Metal Artifact Reduction in CT
@@
$\begin{split}L_{MAR}(\theta)=&\mathbb{E}\left(1+(1-m)\dfrac{max(m\cdot{N})}{max((1-m)\cdot{N})}\right)\\ &\times\|{N_{\theta}}(x_{t},t)-{N}\|_{2}^{2}\end{split}$
@@
E.4 Shadow Removal in Natural Image
@@
The diffused noise N{N} as $(10+\epsilon)({X_{S}}-{x_{0}})/11$ similar to MAR, where the $X_{S}$} is shadow image and $x_{0}$} is non-shadow image. Our learning objective should be consistent with the baseline model ShadowFormer that directly learns the non-shadow image $x_{0}$}. Hence the loss function is: